# Template-Guided 3D Molecular Pose Generation via Flow Matching and Differentiable Optimization

### v1

@@ Abstract
Predicting the 3D conformation of small molecules within protein binding sites is a key challenge in drug design. When a crystallized reference ligand (template) is available, it provides geometric priors that can guide 3D pose prediction. We present a two-stage method for ligand conformation generation guided by such templates. In the first stage, we introduce a molecular alignment approach based on flow-matching to generate 3D coordinates for the ligand, using the template structure as a reference. In the second stage, a differentiable pose optimization procedure refines this conformation based on shape and pharmacophore similarities, internal energy, and, optionally, the protein binding pocket. We evaluate our approach on a new benchmark of ligand pairs co-crystallized with the same target and show that it outperforms standard docking tools and open-access alignment methods, especially in cases involving low similarity to the template or high ligand flexibility.
@@ 1 Introduction
This work introduces Flow Molecular Alignment with Pose Optimization (FMA-PO), a template-guided method for 3D molecular pose generation, which combines LB alignment with structure-aware refinement. FMA-PO employs a Flow Matching (FM) model conditioned on a 3D template molecule. The model generates 3D conformers for a query molecule from its 2D structure, aligning them spatially with the template. The poses are then refined via a coordinate-level differentiable optimization procedure integrating constraints based on shape alignment, pharmacophore similarity, binding pocket complementarity, and internal energy. To evaluate the performance of our method, we introduce AlignDockBench, a new benchmark comprising 369 protein-ligand (PL) template–query pairs. Unlike existing cross-docking benchmarks such as the one proposed by FitDock (Yang et al., 2022)— which focuses on ligand pairs with high chemical similarity—AlignDockBench includes pairs spanning a broader range of similarities. This enables a more challenging and realistic assessment of LB alignment and docking methods.
@@ 3.2 Flow-Matching Molecular Alignment
The FMA module generates 3D conformations of a query ligand conditioned on the 3D structure of a reference ligand. Starting from a random 3D conformation of the query 2D graph, the model progressively denoises it to produce conformations that reproduce the ground-truth pose, spatially aligned with the reference ligand. This alignment process is illustrated in Figure 2, which shows the evolution of the query conformation in blue, from initial noise ($t=0$) to final alignment with the reference structure in pink ($t=1$). Inspired by pocket-guided generative models like DiffDock (Corso et al., 2022), which condition ligand generation on protein pocket geometry, our method instead uses the 3D structure of a template ligand as the conditioning reference.
@@
That is, $\phi_{t}(x)$ represents the point $x$ transported along the vector field $v$ from time 00 to time $t$. To construct a feasible transport trajectory, (Tong et al., 2023) (Albergo et al., 2023) introduces a time-differentiable interpolant:
@@
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\mathbf{x}\sim\rho_{t}(\mathbf{x}_{0},\mathbf{x}_{1})}||\mathbf{v}_{\theta}(t,\mathbf{x})-\mathbf{v}_{t}(\mathbf{x})||^{2},$
@@ 3.3 Pose Optimization
    A pharmacophoric complementarity, using PTS to evaluate the alignment between the query ligand and template pocket pharmacophores. For pocket pharmacophore types, an inversion is applied to model complementary interactions, with illustration provided in Appendix B.2.
@@ 3.4 Benchmark and training dataset construction
AlignDockBench is a curated benchmark designed to evaluate template-based molecular docking and 3D molecular alignment methods. Unlike common docking benchmarks such as PoseBusters (Buttenschoen et al., 2024), which do not include reference ligands, AlignDockBench provides co-crystallized template ligands for each query. This enables the evaluation of methods that leverage known references to guide 3D conformation prediction. The benchmark includes $369$ PL query structures, each associated with a corresponding PL template structure. To construct AlignDockBench, we selected $61$ diverse PL templates from the DUD-E (Mysinger et al., 2012) and DEKOIS (Bauer et al., 2013) datasets, spanning major target classes such as kinases, proteases, nuclear receptors, and GPCRs. For each template, we searched the Protein Data Bank (PDB) for query PL complexes whose binding pockets could be structurally aligned to the pocket template. Binding pockets were defined as all residues within $12$Å of the ligand. Using TM-align (Zhang and Skolnick, ), we rigidly aligned (via translation and rotation) query pocket to the pocket template and retained only those with a backbone Root Mean Square Deviation (RMSD) below $1.2$Å. To ensure sufficient 3D structural similarity between ligands, we further filtered query-template pairs with an STS score greater than $0.5$. The mapping of queries and associated templates is provided in Appendix C.1, along with PL complex preparation details in Appendix C.2. Protein class diversity is summarized in Figure S4. The benchmark is available at Zenodo  [^1] https://anonymous.4open.science/r/AlignDockBench-6756/.
@@ Results.
Table 1 summarizes the performance of each method in terms of mean RMSD, the proportion of molecules with RMSD below $2$Å, and the average runtime per molecule. Both FMA-PO and FMA-PO + outperform competing methods, achieving the lowest mean RMSD and the highest proportion of alignments with RMSD below $2$Å. Examples of poses generated with each method are provided in Appendix D.2.
To assess the effect of template similarity on alignment performance, AlignDockBench was divided into two similarity bins ([0,0.5[[0,0.5[[ 0 , 0.5 [ and $[0.5,1.0]$), based on Tanimoto similarity between the query and template compounds. The similarity was computed using two approaches: first, with standard Morgan fingerprints (Rogers and Hahn, 2010) derived from the complete molecular structures (Figure 3(a)); and second, with Morgan fingerprints computed on the molecules’ Generic Murcko scaffolds (Bemis and Murcko, 1996), which capture the core chemical framework by removing side chains and replacing all atoms with carbons (Figure 3(b)). We further analyzed how molecular complexity affects performance by evaluating results with respect to the number of atoms (Figure 3(c)) and the number of rotatable bonds (Figure 3(d)) in the query molecules. For each bin and each method, we plotted the fraction of molecules achieving RMSD below $2$Å (Figure 3).
@@
All results were obtained under a cross-docking setup, where the pocket structure of the template ligand is used, simulating the realistic scenario in which the query ligand’s pocket structure is unavailable. Additional experiments were conducted under a redocking setup, using the query ligand’s own crystal pocket. Results are provided in the Table S4. We further assessed the quality of generated poses using the PoseBusters test suite (Buttenschoen et al., 2024), with detailed results available in Appendix D.3. An ablation study evaluating the role of the pocket, the importance of the PO stage, and the impact of the number of initial conformations sampled with FMA is also presented in Appendix D.5.
@@
In terms of computational efficiency, FMA-PO achieves a competitive average runtime of $3.66$ seconds per molecule. As expected, FMA-PO + incurs a higher runtime ($27.96$ s per molecule) due to the additional PO steps involving all samples but resulting in improved alignment accuracy. Both FMA-PO and FMA-PO + benefit from GPU  [^3] Experiments were run on a single NVIDIA T4 Tensor Core GPU. acceleration during the FMA stage, whereas conventional methods such as rDock and FitDock were executed using a single CPU core, and Vina was run with four CPU cores.
@@
Method # of Molecules Aligned Mean RMSD (Å ↓) % of Molecules with RMSD $<$ 2Å (↑) Average Runtime (s) FMA-PO 369/369 1.86 ± 1.42 69.38 3.66 $+^{*}$ 369/369 1.62 ± 1.33 77.78 27.96 FitDock 313/369 2.93 ± 3.73 53.67 / 45.53 19.71 LSalign 368/369 2.54 ± 2.00 54.35 / 54.2 5.67 ROSHAMBO 369/369 2.87 ± 2.09 30.35 4.90 rDock 369/369 4.52 ± 3.28 34.42 20.68 Vina 368/369 3.39 ± 2.81 47.28 / 47.15 72.98
@@
Figure 3: Performance of molecular alignment and docking methods on AlignDockBench in a cross-docking scenario. Bar plots (left y-axis) indicate the percentage of molecules with RMSD below $2$Å, while the line plot (right y-axis) shows the number of molecules in each bin. Each figure stratifies the results based on a different structural or chemical property of the query ligand.
@@ 5 Discussion
FMA-PO + is capable of generating 3D poses of candidate compounds that are directly suitable for downstream activity prediction tasks. Indeed, both deep learning-based scoring functions (Shen et al., 2023; Valsson et al., 2025) and physics-based approaches (Greenidge et al., 2013) rely on a 3D representation of the receptor-ligand complex. In addition, using FMA-PO + generated alignments to augment training datasets represents a promising strategy to enhance the performance of predictive or generative models. This aligns with recent efforts such as BindingNet V2 (Zhu et al., 2025), which uses alignment-based data augmentation to boost the performance of deep learning models. In particular, FMA itself could benefit from such augmented data to improve its performance in an iterative training process. Future work could explore conditioning FMA on either a single or a set of reference ligands, along with the protein binding pocket, thereby incorporating more structural context into the pose generation process. Additionally, addressing flexible receptor scenarios, where both the 3D receptor and ligand are predicted, could be interesting. It is worth noting that, compared to non-deep learning methods, FMA-PO remains computationally more intensive, suggesting a potential direction for efficiency improvements.
@@ A.2 Flow-Matching Training and Inference Algorithms
Input: Query 2D graph $G_{query}$, template molecule $\mathcal{M}_{template}$, number of sampling steps $N$, number of samples $K$
@@
Input: Query molecules with 2D gaphs $[G_{query}^{1},...,G_{query}^{K}]$, true coordinates $[C^{1},...,C^{K}]$ and associated template molecules $[\mathcal{M}_{template}^{1},...,\mathcal{M}_{template}^{K}]$, number of epochs $N_{epochs}$, learning rate $\alpha$.
@@ A.3 FMA Molecular Alignment model
FMA (Algorithm 4) takes as input a molecular graph $\mathcal{G}$ representing both query and template ligands, along with their atom coordinates. Positional features $x$ concatenates the ground-truth positions of the template atoms with the query atom coordinates at diffusion step $t$. The model outputs the vector field $v^{query}$ used for query pose prediction. Initially, node and edge features of $\mathcal{G}$ are independently embedded into vectors of dimensions $c_{n}$ and $c_{e}$, respectively, using separate MLPs. All node and edge embeddings are projected into a common feature space with identical dimensionality, thereby transforming the heterogeneous graph into a homogeneous representation. To preserve type-specific information, each node and edge is annotated with its corresponding type attribute. The resulting representation is processed by a stack of $N^{MHA}_{\text{blocks}}$ MHAwithEdgeBias layers [Qiao et al., 2024], jointly updating node and edge embeddings. The processed node embeddings are then passed to the VectorFieldNetwork model (Algorithm 5), which applies a stack of $N^{VFN}_{\text{blocks}}$ layers of PointSetAttentionwithEdgeBias [Qiao et al., 2024], followed by a GatedUpdate layer, adapted from [Qiao et al., 2024] (Algorithm 6). These blocks operate directly on scalar and vector node representations, denoted $f_{s}$ and $f_{v}$ respectively, incorporating positional information $x$ and diffusion time $t$ to produce the final learned vector field $v^{query}$.
@@
_$\>,v\leftarrow\text{{GatedUpdate}}(fs,fv,1)$ $v\in\mathbb{R}^{N_{nodes}\times 3}$
@@ B.1 Shape-based Tanimoto Score
B.2 Pharmacophore-based Tanimoto Score
Pharmacophores of a molecule are 3D arrangements of specific chemical features, such as hydrogen bond donors, hydrogen bond acceptors, hydrophobic groups, and aromatic rings. These features are identified using RDKit. Each pharmacophore feature is positioned at the barycenter of the corresponding atomic group and assigned a specific type (e.g., hydrogen donor, hydrogen acceptor, aromatic, etc.). Pharmacophores are modeled as spherical Gaussian functions with a fixed radius of $1.7$Å, comparable to the van der Waals radius of a carbon atom. The overlapping pharmacophore volume between molecules $A$ and $B$ is given by:

### v2

@@ Abstract
Predicting the 3D conformation of small molecules within protein binding sites is a key challenge in drug design. When a crystallized reference ligand (template) is available, it provides geometric priors that can guide 3D pose prediction. We present a two-stage method for ligand conformation generation guided by such templates. In the first stage, we introduce a molecular alignment approach based on flow-matching to generate 3D coordinates for the ligand, using the template structure as a reference. In the second stage, a differentiable pose optimization procedure refines this conformation based on shape and pharmacophore similarities, internal energy, and, optionally, the protein binding pocket. We introduce a new benchmark of ligand pairs co-crystallized with the same target to evaluate our approach and show that it outperforms standard docking tools and open-access alignment methods, especially in cases involving low similarity to the template or high ligand flexibility.
@@ 1 Introduction
This work introduces Flow Molecular Alignment with Pose Optimization (FMA-PO), a template-guided method for 3D molecular pose generation, which combines LB alignment with structure-aware refinement. FMA-PO employs a Flow Matching (FM) model conditioned on a 3D template molecule. The model generates 3D conformers for a query molecule from its 2D structure, aligning them spatially with the template. The poses are then refined via a coordinate-level differentiable optimization procedure integrating constraints based on shape alignment, pharmacophore similarity, binding pocket complementarity, and internal energy. To evaluate the performance of our method, we introduce AlignDockBench, a new benchmark comprising 369 protein-ligand (PL) template–query pairs. Unlike existing cross-docking benchmarks such as the one proposed by FitDock (Yang et al., 2022)— which focuses on ligand pairs with high chemical similarity—AlignDockBench also includes pairs with lower similarity. This enables a more challenging and realistic assessment of LB alignment and docking methods.
@@ 3.2 Flow-Matching Molecular Alignment
The FMA module generates 3D conformations of a query ligand conditioned on the 3D structure of a reference ligand. Starting from a random 3D conformation of the query 2D graph, the model progressively denoises it to produce conformations that reproduce the ground-truth pose, spatially aligned with the reference ligand. This alignment process is illustrated in Figure 2, which shows the evolution of the query conformation in blue, from initial noise ($t=0$) to final alignment with the reference structure in pink ($t=1$). Inspired by pocket-guided generative models like DiffDock (Corso et al., 2022), which condition ligand generation on protein pocket geometry, our method instead uses the 3D structure of a template ligand as the conditioning reference. An illustration of FMA model pipeline is provided in Figure S1.
(a) $t=0$
(b) $t=0.5$
(c) $t=0.75$
(d) $t=1$
@@
That is, $\phi_{t}(x)$ represents the point $x$ transported along the vector field $v$ from time 0 to time $t$. To construct a feasible transport trajectory, (Tong et al., 2023) (Albergo et al., 2023) introduces a time-differentiable interpolant:
@@
[…]
@@ 3.3 Pose Optimization
    A pharmacophoric complementarity, using PTS to evaluate the alignment between the query ligand and template pocket pharmacophores. For pocket pharmacophore types, an inversion is applied to model complementary interactions. More details are provided in Appendix B.2.
@@ 3.4 Benchmark and training dataset construction
AlignDockBench is a curated benchmark designed to evaluate template-based molecular docking and 3D molecular alignment methods. Unlike common docking benchmarks such as PoseBusters (Buttenschoen et al., 2024), which do not include reference ligands, AlignDockBench provides co-crystallized template ligands for each query. This enables the evaluation of methods that leverage known references to guide 3D conformation prediction. The benchmark includes $369$ PL query structures, each associated with a corresponding PL template structure. To construct AlignDockBench, we selected $61$ diverse PL templates from the DUD-E (Mysinger et al., 2012) and DEKOIS (Bauer et al., 2013) datasets, spanning major target classes such as kinases, proteases, nuclear receptors, and GPCRs. For each template, we searched the Protein Data Bank (PDB) for query PL complexes whose binding pockets could be structurally aligned to the pocket template. Binding pockets were defined as all residues within 1212Å of the ligand. Using TM-align (Zhang and Skolnick, ), we rigidly aligned (via translation and rotation) query pocket to the pocket template and retained only those with a backbone Root Mean Square Deviation (RMSD) below 1.21.2Å. To ensure sufficient 3D structural similarity between ligands, we further filtered query-template pairs with an STS score greater than $0.5$. The mapping of queries and associated templates is provided in Appendix C.1, along with PL complex preparation details in Appendix C.2. Protein class diversity is summarized in Figure S6. The benchmark is available at Zenodo  [^1] https://anonymous.4open.science/r/AlignDockBench-6756/.
@@ Results.
Table 1 summarizes the performance of each method in terms of mean RMSD, the proportion of molecules with RMSD below 22Å, and the average runtime per molecule. Both FMA-PO and FMA-PO + outperform competing methods, achieving the lowest mean RMSD and the highest proportion of alignments with RMSD below 22Å. Examples of poses generated with each method are provided in Figure S13.
To assess the effect of template similarity on alignment performance, AlignDockBench was divided into two similarity bins ($[0,0.5[$ and $[0.5,1.0]$), based on Tanimoto similarity between the query and template compounds. The similarity was computed using two approaches: first, with standard Morgan fingerprints (Rogers and Hahn, 2010) derived from the complete molecular structures (Figure 3(a)); and second, with Morgan fingerprints computed on the molecules’ Generic Murcko scaffolds (Bemis and Murcko, 1996), which capture the core chemical framework by removing side chains and replacing all atoms with carbons (Figure 3(b)). We further analyzed how molecular complexity affects performance by evaluating results with respect to the number of atoms (Figure 3(c)) and the number of rotatable bonds (Figure 3(d)) in the query molecules. For each bin and each method, we plotted the fraction of molecules achieving RMSD below 22Å (Figure 3).
@@
All results were obtained under a cross-docking setup, where the pocket structure of the template ligand is used, simulating the realistic scenario in which the query ligand’s pocket structure is unavailable. Additional experiments were conducted under a redocking setup, using the query ligand’s own crystal pocket. Results are provided in the Table S4. We further assessed the quality of generated poses using the PoseBusters test suite (Buttenschoen et al., 2024), with detailed results available in Appendix D.3. Ablation studies are presented in Appendix D.4. Comparison to Harmonic Flow Stark et al. (2024) is provided in D.2.3.
@@
In terms of computational efficiency, FMA-PO achieves a competitive average runtime of $3.66$ seconds per molecule. As expected, FMA-PO + incurs a higher runtime ($27.96$ s per molecule) due to the additional PO steps involving all samples but resulting in improved alignment accuracy. Both FMA-PO and FMA-PO + benefit from GPU  [^3] Experiments were run on a single NVIDIA T4 Tensor Core GPU. acceleration during the FMA stage, whereas conventional methods such as rDock and FitDock were executed using a single CPU core, and Vina was run with four CPU cores. More details are provided in D.2.4.
@@
[…]
@@
Figure 3: Performance of molecular alignment and docking methods on AlignDockBench in a cross-docking scenario. Bar plots (left y-axis) indicate the percentage of molecules with RMSD below 22Å, while the line plot (right y-axis) shows the number of molecules in each bin. Each figure stratifies the results based on a different structural or chemical property of the query ligand.
@@ 5 Discussion
FMA-PO + is capable of generating 3D poses of candidate compounds that are directly suitable for downstream activity prediction tasks. Indeed, both deep learning-based scoring functions (Shen et al., 2023; Valsson et al., 2025) and physics-based approaches (Greenidge et al., 2013) rely on a 3D representation of the receptor-ligand complex. In addition, using FMA-PO + generated alignments to augment training datasets represents a promising strategy to enhance the performance of predictive or generative models. This aligns with recent efforts such as BindingNet V2 (Zhu et al., 2025), which uses alignment-based data augmentation to boost the performance of deep learning models. In particular, FMA itself could benefit from such augmented data to improve its performance in an iterative training process. Future work could explore conditioning FMA on either a single or a set of reference ligands, along with the protein binding pocket, thereby incorporating more structural context into the pose generation process Guan et al. (2025). Additionally, addressing flexible receptor scenarios, where both the 3D receptor and ligand are predicted, could be interesting. It is worth noting that, compared to non-deep learning methods, FMA-PO remains computationally more intensive, suggesting a potential direction for efficiency improvements.
@@ 6 Conclusion
Acknowledgment
We would like to express our sincere gratitude to the Iktos team for their exceptional contributions to this project. The authors would like to thank IKTOS for supporting this study. We are also grateful to the reviewers for their valuable comments and suggestions, which have improved the manuscript. We further thank Ennys Gheyouche, Maoussi Lhuillier-Akakpo, and Juan Sanz García for their feedback and review of an early draft of this work.
@@ A.2 Flow-Matching Training and Inference Algorithms
Input: Query 2D graph $G_{query}$, template molecule $\mathcal{M}_{template}$, number of sampling steps NN, number of samples $K$
@@
Input: Query molecules with 2D graphs $G_{query}^{1},...,G_{query}^{K}]$, true coordinates $C^{1},...,C^{K}]$ and associated template molecules $\mathcal{M}_{template}^{1},...,\mathcal{M}_{template}^{K}]$, number of epochs $N_{epochs}$, learning rate $\alpha$.
@@ A.3 FMA Molecular Alignment model
FMA (Algorithm 4) takes as input a molecular graph $\mathcal{G}$ representing both query and template ligands, along with their atom coordinates. Positional features $x$ concatenates the ground-truth positions of the template atoms with the query atom coordinates at diffusion step $t$. The model outputs the vector field $v^{query}$ used for query pose prediction. Initially, node and edge features of $\mathcal{G}$ are independently embedded into vectors of dimensions $c_{n}$ and $c_{e}$, respectively, using separate MLPs. All node and edge embeddings are projected into a common feature space with identical dimensionality, thereby transforming the heterogeneous graph into a homogeneous representation. To preserve type-specific information, each node and edge is annotated with its corresponding type attribute. The resulting representation is processed by a stack of $N^{MHA}_{\text{blocks}}$ MHAwithEdgeBias layers (Algorithm S3 in [Qiao et al., 2024]), jointly updating node and edge embeddings. The processed node embeddings are then passed to the VectorFieldNetwork model (Algorithm 5), which applies a stack of $N^{VFN}_{\text{blocks}}$ layers of PointSetAttentionwithEdgeBias (Algorithm S8 in [Qiao et al., 2024]), followed by a GatedUpdate layer (Algorithm 6), adapted from [Qiao et al., 2024]. These blocks operate directly on scalar and vector node representations, denoted $f_{s}$ and $f_{v}$ respectively, incorporating positional information $x$ and diffusion time $t$ to produce the final learned vector field $v^{query}$.
@@
$v\leftarrow\text{{GatedUpdate}}(fs,fv,1)$ $v\in\mathbb{R}^{N_{nodes}\times 3}$
@@ B.1 Shape-based Tanimoto Score
B.2 Pharmacophore scoring
Pharmacophore-based Tanimoto Score. Pharmacophores of a molecule are 3D arrangements of specific chemical features, such as hydrogen bond donors, hydrogen bond acceptors, hydrophobic groups, and aromatic rings. These features are identified using RDKit. Each pharmacophore feature is positioned at the barycenter of the corresponding atomic group and assigned a specific type (e.g., hydrogen donor, hydrogen acceptor, aromatic, etc.). Pharmacophores are modeled as spherical Gaussian functions with a fixed radius of 1.71.7Å, comparable to the van der Waals radius of a carbon atom. The overlapping pharmacophore volume between molecules $A$ and $B$ is given by:
@@ B.2 Pharmacophore scoring
Pharmacophoric complementarity. The concept of pharmacophoric complementarity in our method stems from Gaussian-based shape similarity modeling. Specifically, once a Gaussian sphere is labeled with a pharmacophoric feature (e.g., hydrogen bond donor (HBD), hydrophobic, etc.), it becomes possible to compute a pharmacophoric similarity between two ligands based on the overlap of these labeled Gaussian volumes.
When introducing the protein pocket into the PO process, we extended this idea to model complementarity between ligand and pocket by maximizing the overlap between complementary pharmacophoric Gaussians. For instance, Hydrogen Bond Donor (HBD) and Hydrogen Bond Acceptor (HBA) were defined as complementary, as well as cationic and anionic features (Figure S4). During the PO, the $\mathcal{S}_{\mathrm{PTS}}$ term of the objective function increases the overlap between such complementary features from the ligand and the pocket, whereas the shape $\mathcal{S}_{\mathrm{STS}}$ term prevents steric clashes by penalizing excessive overlap between atoms.
Regarding hydrogen bonding in particular, we model the HBD pharmacophoric center as a Gaussian sphere located at the position of the donor hydrogen atom. If this hydrogen atom (and thus its associated Gaussian sphere) is positioned near an HBA-labeled Gaussian from the pocket, the overlap between these spheres is maximized, effectively favoring geometries consistent with hydrogen bond formation. The geometry naturally tends toward a favorable $\mathrm{X}{-}\mathrm{H}\cdot\cdot\cdot\mathrm{X}$ angle close to $180^{\circ}$, since this configuration maximizes Gaussian overlap. Simultaneously, the internal energy term in the PO loss encourages the donor group (e.g., $\mathrm{C}{-}\mathrm{X}{-}\mathrm{H}$) to remain in energetically favorable conformations.
@@
As an illustrative example, consider the protein–ligand complex (PDB: 1NJS) shown in Fig. S5. The ligand KEU (sequence number 510, chain A) contains two hydroxyl groups, C5-OA1 and C5-OA2, both positioned near the carboxylate group of ASP 144 (chain A):
  • The hydroxyl oxygen OA1 belongs to a donor group (C5-OA1-H1) and lies near the acceptor OD1 of ASP 144.
  • The second hydroxyl oxygen OA2 (in C5-OA2-H2) is close to the second acceptor OD2.
At the beginning of PO, hydrogen atoms H1 and H2 are added by RDKit without considering the protein pocket. However, once optimization begins, complementary HBD Gaussians are assigned to these hydrogens, and their overlap with HBA Gaussians on OD1 and OD2 is maximized.
Figure S5: Illustration of iterative optimization of hydrogen bonding interactions in a protein–ligand complex. 3D structure of the human GAR Tfase (carbons in cyan) in complex with the hydrolyzed form of 10-trifluoroacetyl-5,10-dideaza-acyclic-5,6,7,8-tetrahydrofolic acid Zhang et al. [2003] (carbons in grey). (A) The protein–ligand binding site shows two HBD (Donor1, Donor2) and two HBA (Acceptor1, Acceptor2). (B) Time evolution of donor–acceptor distances ($\mathrm{H1\text{--}Acceptor1}$, $\mathrm{H2\text{--}Acceptor2}$) over the optimization iterations. (C) Time evolution of hydrogen-bond angles ($\mathrm{Donor1\text{--}H1\text{--}Acceptor1}$, $\mathrm{Donor2\text{--}H2\text{--}Acceptor2}$). (D) Representative ligand conformations at selected iterations (20, 40, 70, 113), illustrating the progressive stabilization of hydrogen bonds in the binding pocket. (E) Evolution of the angles between the carbon atoms bound to hydrogen-bond donors and their respective hydrogens ($\mathrm{C\text{--}H_{Donor1}\text{--}H1}$, $\mathrm{C\text{--}H_{Donor2}\text{--}H2}$). (F) GAFF2 energy profile across iterations, showing convergence from initially high-energy conformations toward a stabilized minimum.